A border patrol helicopter set out to catch a drug smuggler on a speed boat. The speed
boat is travelling south of manila going 27m/sec and left 2 hours ago. If the helicopter
leaves now and travel 67 m/sec. how long will it take to catch the smuggler.
1
Expert's answer
2021-07-23T05:18:17-0400

The initial distance between boat and helicopter

"d=v_b*\Delta t=(27\:\rm m/s)*2*3600\: s=194400\: m"

The time needed for a border patrol to catch the smuggler

"t=\frac{d}{v_{rel}}=\frac{d}{v_h-v_b}\\
=\frac{194400\:\rm m}{(67-27)\:\rm m/s}=4860\: \rm s=81\: min=1.35 \: hr"